OpenStudy (anonymous):

The sequence is geometric. What is the next term? 3125, 625, 125, 25, ... 20 15 5

OpenStudy (hoblos):

25/125= 1/5 so the ratio is 1/5 25 * (1/5) =5

OpenStudy (anonymous):

this is a G.s ,therfore the diffrence between the two terms is Tn+1\r=tn so the r here = 0.2, therfore last term is 25 ,next term = 25*0.2=5